OpenStudy (anonymous):

Find y' and y''. y = x^2 ln(9x) y' = y''=

OpenStudy (anonymous):

Product rule, Power rule, Chain rule. Those are the rules you need to use. @TachiHere

OpenStudy (anonymous):

I'll do the first one for you:\[\bf y'=[x^2\ln(9x)]'=(x^2)'\ln(9x)+x^2(\ln(9x))'=2xln(9x)+x^2\left( \frac{ 1 }{ \cancel9x } \right)\cancel9\]\[\ \ \ \bf = 2xln(9x)+x\]Also note that the derivative of ln(u) where 'u' is a function of 'x' is:\[\bf \ln(u)=\frac{ 1 }{ u }*\frac{ du }{ dx }\]Now can you find y''? @TachiHere

OpenStudy (cwrw238):

2x ln(9x) is the product of 2 functions so use product rule for tis part